OpenStudy (anonymous):

Medicine Hat is 100km east of Lethbridge. roy left his home in Medicine hat and drove east at 90 km/hr a) write the partial variation equation thay relates Roys driving distance from Lethbridge to his driving time . b) after driving for 2.5h how far is roy from lethbridge?

OpenStudy (anonymous):

Edit: Sign mistake a) distance(time) = 100 + 90 * time Distance in km, time in hour b) distance(2.5) = 100 + 90 * 2.5 = 325 Roy is 325 km away from lethbridge